OpenStudy (anonymous):

Simplify (x + 3)(x2 + 4x – 2)

OpenStudy (anonymous):

= (x + 3)(x² + 4x - 2) = x³ + 4x² - 2x + 3x² + 12x - 6 = x³ + 7x² + 10x - 6 x³ + 7x² + 10x - 6

OpenStudy (anonymous):

\[\Large x^3+4x^2-2x+3x^2+12x-6\] now put terms together ... and you'll have ..??

OpenStudy (anonymous):

@Keyonn that's correct, well done !

OpenStudy (anonymous):

so did i get the wrong answer ? x³ + 7x² + 10x - 6